# NutrifyAI: An AI-Powered System for Real-Time Food Detection, Nutritional Analysis, and Personalized Meal Recommendations

### v1

@@ VI Conclusion
In summary, NutrifyAI demonstrates the potential of integrating computer vision models with nutritional APIs to create an effective tool for food recognition and nutritional analysis. While the system achieved promising results, several areas remain for further exploration and enhancement. Future research should focus on expanding the diversity of training datasets to improve the model’s ability to recognize a wider array of food items. Additionally, refining the model’s confidence calibration and incorporating advanced evaluation metrics, such as mAP, will be crucial for achieving greater reliability. Developing mechanisms for user feedback will enable the system to adapt to real-world usage, offering more personalized and accurate meal recommendations. As the field of AI-powered nutrition continues to evolve, there is significant potential for this technology to be applied in various domains, including healthcare and personalized diet planning, making the pursuit of these improvements highly worthwhile.
Acknowledgment
This project would not have been possible without the support and guidance of Junyao Chen and Professor Zhengyuan Zhou from the NYU Stern School of Business, Department of Technology, Operations, and Statistics. Their mentorship and insights were crucial in shaping the direction of this research. Furthermore, special thanks to the Winston Data Foundation for their financial support through their merit scholarship for the NYU GSTEM Program.

### v2

@@ V Conclusion
NutrifyAI demonstrates the potential of integrating computer vision models with nutritional APIs for food recognition and nutritional analysis. Despite its promising results, the system could benefit from further improvements. Future efforts should aim to diversify training datasets to enhance food item recognition and refine confidence calibration to increase reliability. Implementing user feedback mechanisms will enable the system to adapt to real-world usage, offering more personalized and accurate meal recommendations. As the field of AI-powered nutrition continues to evolve, there is significant potential for this technology to be applied in various domains, including healthcare and personalized diet planning, making the pursuit of these improvements highly worthwhile.